# Unlocking Adversarial Suffix Optimization Without Affirmative Phrases: Efficient Black-box Jailbreaking via LLM as Optimizer

### v1

@@ 3.4 Algorithm
The detailed algorithm is illustrated in Algorithm 1. This method constitutes an iterative optimization framework, allowing for up to $K$ rounds of suffix optimization for a given malicious query (in Line 3). As previously stated, we provide the LLM with historical references for self-reflection, commencing with a sampling-based selection on the reference history list (in Line 4). It is notable that a maximum number of reference history pairs is designated as $r$. Thus, in the beginning, when the history list is empty, the reference prompt for that round is omitted. Furthermore, the malicious query and historical references are integrated into an attacker template to construct the task prompt (in Line 5). By feeding this integrated prompt to the target LLM, candidate suffixes are generated (in Line 6). To enhance optimization efficiency, a batch generation strategy is employed to produce multiple candidate suffixes simultaneously. Each generated suffix is evaluated by a scorer to assess its efficacy as an attack and to determine whether the jailbreak is successful (in Line 9-13). If successful, the generated jailbreaking prompt is returned (in Line 10-11). Otherwise, the current suffix and its score are added to the history list, and the process moves to the next optimization round (in Line 13) until a jailbreak is successful or the iteration limit is reached.
@@ 4.1 Experimental Setups
Our method is implemented with Python 3.8 and PyTorch 2.1. All experiments are conducted on a Ubuntu 20.04 server equipped with four NVIDIA A800 GPUs.
Baselines. To assess the performance of ECLIPSE, we compared it against three baselines: the optimization-based method GCG [11] and two template-based methods, DeepInception [10], GPTFUZZER [9] and PAIR [12]. ECLIPSE was configured to generate suffixes with temperature as 1.2 and conduct up to $k$ = 50 rounds of iterations, producing $b$ = 8 candidate answers in one batch, with selection $r$ = 10 historical references, totaling up to 400 generate prompts. GCG operated with its default batch size of 512 across 500 rounds, totaling 256,000 trials. DeepInception, using manually crafted templates, required only one trial per query. GPTFUZZER, similar to ECLIPSE, conducted 400 trials per query. All other settings in baselines were kept at their original defaults. Additionally, we factored in the influence of specific LLM tokens, such as [INST], known to affect jailbreaking effectiveness, by incorporating the “[INST] input [/INST]” pattern in our setups, following insights from recent studies [31].
Models. Our evaluation covered three popular open-source LLMs: LLaMA2-7B-Chat [32], Vicuna-7B [33] and Falcon-7B-Instruct [34]. We also involved the commercial GPT-3.5-Turbo API [1]. Each LLM was allowed to generate up to 256 tokens for every malicious query. For ease of presentation, hereafter we will refer to LLaMA2-7B-Chat, Vicuna-7B, Falcon-7B-Instruct, and GPT-3.5-Turbo as LLaMA2, Vicuna, Falcon, and GPT-3.5, respectively.
Datasets. We utilized the AdvBench dataset, a widely recognized benchmark in jailbreaking research and employed in GCG [11], comprising 520 malicious queries with corresponding affirmative phrases spanning various categories like profanity, graphic content, and cybercrime. For our experimental analysis, we extracted 100 malicious questions from this dataset. To enable a comparison with template-based DeepInception, we initially sampled 100 questions from their dataset, identifying 66 unique malicious queries after deduplication. This dual dataset approach was crucial to account for the distinct content and methodologies inherent to GCG and DeepInception, with GCG focusing on affirmative phrases and DeepInception on crafted templates.
@@
Metrics. To thoroughly evaluate ECLIPSE from multiple perspectives, we utilized the following metrics: Attack Success Rate (ASR), Query Rounds (QR), Query Numbers (QN), Overhead (QH), and Perplexity (PPL). The attack success rate assesses the effectiveness by measuring the proportion of malicious queries from a set of 100 AdvBench questions for which ECLIPSE successfully identifies adversarial suffixes leading to a jailbreak. We employed a dual assurance strategy for assessing the effectiveness of attacks, with both refusal-to-answer matching [11] and automated classifiers [9] simultaneously. We were aware of the recent practice of using the LLM-as-a-Judge approach [12]. Due to budget constraints, we opted for GPT-3.5 instead of GPT-4 for a trial. However, we encountered frequent false negatives with GPT-3.5, which led us to abandon this choice. The query rounds, query numbers, and overhead reflect the efficiency, indicating the number of optimization rounds and queries required and the computational time cost measured in seconds, respectively. When comparing with GCG, we compared the QR since GCG utilized a larger batch size. When comparing with templated-based methods, we directly compared the QN. Considering the concern that the sequences generated by token-based optimization may be strange and poorly stealthy, we measured the standard sentence perplexity (PPL) based on GPT-2, with a lower value representing more naturalness.
@@ 4.2 Performance Evaluation
Moreover, ECLIPSE marked considerable advancements in efficiency, significantly reducing both the number of query rounds (QR) and overhead (OH). For instance, on LLaMA2, QR was reduced from 178.5 to 16.81 and OH from 1557.39 seconds to 173.45 seconds, demonstrating a more than tenfold improvement in speed and efficiency. Similar enhancements were noted with Vicuna, emphasizing ECLIPSE’s capability to optimize attack strategies effectively. For Falcon, ECLIPSE not only achieved high effectiveness with an ASR of 0.98 but also showcased remarkable efficiency, requiring only 3.95 optimization rounds and 28.67 seconds—significantly better than GCG’s performance, which only achieved an ASR of 0.21 ECLIPSE-generated jailbreaking prompts also exhibited significantly lower perplexity (PPL) than GCG’s prompts for LLaMA2 and Falcon, indicating enhanced naturalness. However, for Vicuna, where GCG’s prompts exhibited very low PPL, this was likely due to Vicuna’s limited alignment, where the default initialized suffix may induce malicious responses.
Comparison with Template-based Methods.
We conducted evaluations on the LLaMA2 and GPT-3.5 models, where our method not only achieved results comparable to template-based methods but also demonstrated higher optimization efficiency, as detailed in  Table 3. On the LLaMA2, ECLIPSE achieved an ASR of 0.74, surpassing both the 0.52 ASR of GPTFUZZER and the 0.53 ASR of PAIR. For GPT-3.5, while PAIR reached a perfect ASR of 1.0, similar to GPTFUZZER, ECLIPSE was slightly less effective with an ASR of 0.94. Despite this, ECLIPSE showed significant advantages in terms of efficiency across both models: the number of queries was reduced by an average of 45%, and overhead was cut by an average of 83%. On GPT-3.5, we observed a particularly poor performance from DeepInception, likely due to the model’s strong adherence to instructions, which led it to assume the role of the template excessively, shifting focus towards storytelling rather than addressing the original malicious queries. ECLIPSE outperformed PAIR in terms of efficiency, despite with more QR. This is because our task template was simpler and more direct, whereas PAIR’s approach consumed more tokens, leading to a slower process despite its effectiveness.
@@ 4.4 Ablation Studies
Removing the historical references. In our study, we used historical suffixes and scores to boost the LLM’s optimization capabilities. We investigated whether ECLIPSE could still produce effective jailbreaking suffixes without these historical references by conducting an ablation study, the results of which are presented in Table 4. The findings revealed a significant decline in performance when historical references were removed. For example, the ASR for the LLaMA2 model plummeted from 75% to 35%. Additionally, the absence of historical data led to an increase in the number of query rounds and overhead across all models. For instance, query rounds for Vicuna increased from 3.41 to 9.23, and overhead for GPT-3.5 went up from 93.87 seconds to 131.22 seconds. These results underscore the importance of historical data in maintaining the efficiency and effectiveness of ECLIPSE in generating jailbreaking suffixes.
Removing the hidden space features (HSF) in task prompting. We have mentioned that we prompted the LLM to act as an attacker by identifying suffixes that could influence the hidden space features (HSF) of a given query. Here, we delved into the impact of this component.  Table 4 illustrated the effects of removing this instruction from the task prompts; the effectiveness of ECLIPSE on the LLaMA2 model was notably compromised, with the ASR decreasing from 0.75 to 0.59, a 21% reduction. And for GPT-3.5, the ASR also dropped from 0.97 to 0.94. On other models, while the ASR was not significantly affected, there was a slight decrease in attack efficiency, with both QR and OH experiencing marginal increases. Detailed task prompts can be found in the Appendix.
@@
5 Discussions
Limitations. Similar to existing template-based methods [8, 9, 10, 12], our method also needs to leverage some of the instruct-following capability of LLMs to perform the generation and optimization. However, our method only needs to have the LLM generate a suffix, which is easier than template-based methods that utilize LLMs for complex tasks like sentence rewriting and story design. We have assessed the performance of our methods across multiple LLMs in the evaluation section.
Ethics. Research on jailbreaking LLMs raises some ethical concerns [35, 36, 31, 37, 38, 39, 40, 41], e.g., generating harmful and illegal content. However, jailbreaking methods can also serve as effective red-teaming tools to examine and evaluate the current safety alignment of LLMs. We believe that our method will contribute to enhancing the robustness and safety of LLMs.
@@ 6 Conclusion
Appendix A Appendix
A.1 Task Prompts Involved
This section presents the specific task prompts utilized in this paper. Additionally, it includes the prompts used to observe the impact on jailbreak performance with the removal of content related to hidden space features from the prompts.
@@ A.1 Task Prompts Involved
A.2 A Case Study of Numeric Suffixes
During our early experiments, we discover an intriguing phenomenon. Without any specific hints, the attacker LLM tends to actively generate some numeric suffixes. These numbers, when appending directly to the given malicious query and input into the LLM, significantly increase the likelihood of the LLM to continue writing extensive and diverse harmful content. Note that we explore the LLM itself here, i.e., the query is entered without using any other dialog template. However, when we replace these numbers with random numbers, the success rate of inducing harmful content drastically decreases. This observation suggests that LLMs may indeed possess and share certain special knowledge or memories, which can be exploited to induce harmful behaviors.
@@ A.2 A Case Study of Numeric Suffixes
We make two further observations, with some empirical experiments on LLaMA2-7B-Chat. Firstly, we find that a numerical suffix may be able to activate certain specific memories in LLM. For example, as shown in  Table 5, adding a suffix “7.3” to Case 1 induces LLM to generate a sentence similar to a log mark, and adding a suffix “6826.2” to Case 2 induces LLM to generate the address of a complementary pdf. Although we are currently unable to verify the authenticity of these logs or addresses, this responds to the possibility that LLM has seen similar content in the training data. Secondly, we try to analyze why adding a number causes LLM to behave differently. More specifically, we randomly selected some malicious queries, enumerated tens of thousands of numeric suffixes for each query, and verified whether they induced harmful content. We visualize the embeddings of the prompts that successfully induced harmful content along with those that failed, wich PCA. Figure 4 shows a clear difference in the distribution of embeddings for failed prompts (red dots) and successful prompts (blue dots), which enlightens us that some numerical suffixes greatly perturb the hidden space features of the original malicious queries, thus destroying the robustness of the safety alignment.
A.3 A Study on Longitudinal Versions of GPT-3.5-Turbo
@@
enhancing linguistic capabilities may have compromised some aspects of robustness and security [42]. To explore this, we conduct a study on three versions of the GPT-3.5-Turbo model, specifically the 2023-0613, 2023-1106, and 2024-0125 releases, with ECLIPSE. As illustrated in  Figure 5, our results also corroborate this viewpoint. With updates across three versions of GPT-3.5-Turbo, the ASR increases progressively from 0.37 to 0.81, and then to 0.97, indicating a growing likelihood of producing harmful content.

### v2

@@ 3.4 Algorithm
The detailed algorithm is illustrated in Algorithm 1. This iterative optimization framework allows up to $K$ rounds of suffix optimization for a given malicious query (Line 3). As previously stated, we provide the LLM with historical references for self-reflection by sampling from the reference history list (Line 4), with a maximum of $r$ reference pairs. Initially, when the history list is empty, the reference prompt is omitted. The malicious query and historical references are then integrated into an attacker template to construct the task prompt (Line 5). This prompt is fed to the target LLM to generate candidate suffixes (Line 6). To enhance optimization efficiency, a batch generation strategy is employed to produce multiple candidate suffixes simultaneously. Each generated suffix is evaluated by a scorer to assess its attack efficacy and determine whether the jailbreak is successful (Lines 9–13). If successful, the generated jailbreak prompt is returned (Lines 10–11); otherwise, the current suffix and its score are added to the history list, and the process advances to the next optimization round (Line 13) until a successful jailbreak is found or the iteration limit is reached.
@@ 4.1 Experimental Setups
Our method is implemented with Python 3.8. All experiments are conducted on a Ubuntu 20.04 server with four NVIDIA A800 GPUs.
Baselines. We evaluated ECLIPSE against state-of-the-art jailbreaking methods, including the optimization-based GCG Zou et al. (2023) and template-based approaches (DeepInception Li et al. (2023), GPTFUZZER Yu et al. (2023), and PAIR Chao et al. (2023)). For fair comparison, we carefully configured each method while considering their distinct characteristics. For ECLIPSE, we employed the following configuration: temperature of 1.2, maximum iteration rounds $k$ = 50, batch size $b$ = 8 for candidate answers, and $r$ = 10 historical references, culminating in 400 generated prompts per query. For baselines, we followed established practices: GCG was executed with its standard configuration (batch size 512, 500 rounds), resulting in 256,000 prompts per query - notably more resources than other methods yet achieving lower attack success rates as shown in § 4.2. For template-based approaches, we maintained consistency by setting a 400-prompt limit for both GPTFUZZER and PAIR, with PAIR sharing similar parameters to ECLIPSE for iterations and parallel prompt generation. DeepInception, being template-based with fixed manual patterns, requires only one prompt per query and thus serves primarily for effectiveness rather than efficiency comparison. All other baseline parameters remained at their default values as reported in their respective papers. Additionally, we factored in the influence of specific LLM tokens, such as [INST], known to affect jailbreaking effectiveness, by incorporating the “[INST] input [/INST]” pattern in our setups, following insights from recent studies Xu et al. (2024).
Models. Our evaluation covered three popular open-source LLMs: LLaMA2-7B-Chat Touvron et al. (2023), Vicuna-7B Chiang et al. (2023) and Falcon-7B-Instruct Almazrouei et al. (2023). We also involved the commercial GPT-3.5-Turbo API OpenAI (2023). Each LLM was allowed to generate up to 256 tokens for every malicious query. For ease of presentation, hereafter we will refer to LLaMA2-7B-Chat, Vicuna-7B, Falcon-7B-Instruct, and GPT-3.5-Turbo as LLaMA2, Vicuna, Falcon, and GPT-3.5, respectively.
Datasets. We utilized the AdvBench dataset, a widely recognized benchmark in jailbreaking research and employed in GCG Zou et al. (2023), comprising 520 malicious queries with corresponding affirmative phrases spanning various categories like profanity, graphic content, and cybercrime. For our experimental analysis, we extracted 100 malicious questions from this dataset. To enable a comparison with template-based DeepInception, we initially sampled 100 questions from their dataset, identifying 66 unique malicious queries after deduplication. This dual dataset approach was crucial to account for the distinct content and methodologies inherent to GCG and DeepInception, with GCG focusing on affirmative phrases and DeepInception on crafted templates.
@@
Metrics. To thoroughly evaluate ECLIPSE from multiple perspectives, we utilized the following metrics: Attack Success Rate (ASR), Query Rounds (QR), Query Numbers (QN), Overhead (QH), and Perplexity (PPL). The attack success rate assesses the effectiveness by measuring the proportion of malicious queries from a set of 100 AdvBench questions for which ECLIPSE successfully identifies adversarial suffixes leading to a jailbreak. We employed a dual assurance strategy for assessing the effectiveness of attacks, with both refusal-to-answer matching Zou et al. (2023) and automated classifiers Yu et al. (2023) simultaneously. We were aware of the recent practice of using the LLM-as-a-Judge approach Chao et al. (2023). Due to budget constraints, we opted for GPT-3.5 instead of GPT-4 for a trial. However, we encountered frequent false negatives with GPT-3.5, which led us to abandon this choice. The query rounds, query numbers, and overhead reflect the efficiency, indicating the number of optimization rounds and queries required and the computational time cost measured in seconds, respectively. When comparing with GCG, we compared the QR since GCG utilized a larger batch size. When comparing with templated-based methods, we directly compared the QN. Addressing concerns about the naturalness of generated jailbreak prompts, we assess their perplexity (PPL) using GPT-2. Lower PPL values indicate greater fluency and naturalness, mitigating detectability.
@@ 4.2 Performance Evaluation
Moreover, ECLIPSE marked considerable advancements in efficiency, reducing both the number of query rounds (QR) and overhead (OH). For instance, on LLaMA2, QR was reduced from 178.5 to 16.81 and OH from 1557.39 seconds to 173.45 seconds, demonstrating a more than tenfold improvement. Similar enhancements were noted with Vicuna, emphasizing ECLIPSE’s capability to optimize attack strategies effectively. For Falcon, ECLIPSE not only achieved high effectiveness with an ASR of 0.98 but also showcased remarkable efficiency, requiring only 3.95 optimization rounds and 28.67 seconds—significantly better than GCG, which only achieved an ASR of 0.21 our generated prompts achieved lower perplexity (PPL) than GCG’s outputs on LLaMA2 and Falcon, demonstrating superior linguistic naturalness. The exception was Vicuna, where GCG’s low PPL likely stemmed from its weak alignment, allowing even default suffixes to trigger harmful responses.
Comparison with Template-based Methods. We conducted evaluations on the LLaMA2 and GPT-3.5, where our method not only achieved results comparable to template-based methods but also demonstrated higher optimization efficiency, as detailed in  Table 3. On the LLaMA2, ECLIPSE achieved an ASR of 0.74, surpassing both the 0.52 ASR of GPTFUZZER and the 0.53 ASR of PAIR. For GPT-3.5, while PAIR reached a perfect ASR of 1.0, similar to GPTFUZZER, ECLIPSE was slightly less effective with an ASR of 0.94. Despite this, ECLIPSE showed significant advantages in terms of efficiency across both models: the number of queries was reduced by an average of 45%, and overhead was cut by an average of 83%. On GPT-3.5, we observed a particularly poor performance from DeepInception, likely due to the model’s strong adherence to instructions, which led it to assume the role of the template excessively, shifting focus towards storytelling rather than addressing the original malicious queries. ECLIPSE outperformed PAIR in terms of efficiency, despite with more QR. This is because our task template was simpler and more direct, whereas PAIR’s approach consumed more tokens, leading to a slower process despite its effectiveness.
@@ 4.3 Exploring the Transferability of Attacker
4.4 Potential Defenses
In addition to evaluating stealthiness through perplexity scores where ECLIPSE-generated jailbreak suffixes demonstrated superior performance, we further assessed our method against popular defense mechanisms. Specifically, we conducted comprehensive evaluations using several industry-standard content safety checkers: Salesforce Checker Salesforce (2024), LlamaGuard Team (2024), and Azure Checker Azure (2024). Table 5 summarizes their detection rates for different models. For the LLaMA2 model, the detection rates are as follows: Salesforce Checker detects approximately $6\%$, LlamaGuard detects about $38\%$, and Azure Checker detects roughly $46\%$ of the harmful outputs. In contrast, for the GPT-3.5 model, the detection rates are higher, with $11\%$, $72\%$, and $69\%$ detected by Salesforce Checker, LlamaGuard, and Azure Checker, respectively. This indicates that GPT-3.5 tends to generate harmful responses that are more readily detected by these safety mechanisms, potentially due to its ability to produce higher-quality outputs that inadvertently reveal harmful content.
Table 5: Detection rates with various content safety checkers.
Model Salesforce LlamaGuard Azure
LLaMA2-7B-Chat $6\%$ $38\%$ $46\%$
GPT-3.5-Turbo $11\%$ $72\%$ $69\%$
@@ 4.5 Ablation Studies
Removing the historical references. In our study, we used historical suffixes and scores to boost the LLM’s optimization capabilities. We investigated whether ECLIPSE could still produce effective jailbreaking suffixes without these historical references. The results are presented in Table 4. The findings revealed a significant decline in performance when historical references were removed. For example, the ASR for the LLaMA2 model plummeted from 75% to 35%. Additionally, the absence of historical data led to an increase in the number of query rounds and overhead across all models. For instance, query rounds for Vicuna increased from 3.41 to 9.23, and overhead for GPT-3.5 went up from 93.87 seconds to 131.22 seconds. These results underscore the importance of historical data in maintaining the efficiency and effectiveness of ECLIPSE.
Removing the hidden space features (HSF) in task prompting. We have mentioned that we prompted the LLM to act as an attacker by identifying suffixes that could influence the hidden space features (HSF) of a given query. Here, we delved into the impact of this component.  Table 4 illustrated the effects of removing this instruction from the task prompts; the effectiveness of ECLIPSE on the LLaMA2 model was notably compromised, with the ASR decreasing from 0.75 to 0.59, a 21% reduction. And for GPT-3.5, the ASR also dropped from 0.97 to 0.94. On other models, while the ASR was not significantly affected, there was a slight decrease in attack efficiency, with both QR and OH experiencing marginal increases. Detailed task prompts can be found in the Appendix A.
@@ 5 Conclusion
Ethical Considerations and Limitations
Ethics. Research on jailbreaking LLMs raises some ethical concerns Zhang et al. (2023); Wei et al. (2024); Xu et al. (2024); Kumar et al. (2023); Ji et al. (2024); Tian et al. (2023); Zheng et al. (2024); Ren et al. (2024), e.g., generating harmful and illegal content. However, jailbreaking methods can also serve as effective red-teaming tools to examine and evaluate the current safety alignment of LLMs. We believe that our method will contribute to enhancing the robustness and safety of LLMs.
Limitations. Similar to existing template-based methods Deng et al. (2023); Yu et al. (2023); Li et al. (2023); Chao et al. (2023), our method also needs to leverage some of the instruct-following capability of LLMs to perform the generation and optimization. However, our method only needs to have the LLM generate a suffix, which is easier than template-based methods that utilize LLMs for complex tasks like sentence rewriting and story design. We have assessed the performance of our methods across multiple LLMs in the evaluation section.
Acknowledgments
Weipeng Jiang, Zhengyu Zhao, and Chao Shen are supported by the National Key Research and Development Program of China (2023YFE0209800), the National Natural Science Foundation of China (U24B20185, T2442014, 62161160337, 62132011, U21B2018), the Shaanxi Province Key Industry Innovation Program (2023-ZDLGY-38, 2021ZDLGY01-02).
Appendix A Task Prompts Involved
Here, we list in detail the prompts templates used in the experiments. The first two are standard templates for employing the attacker LLM by ECLIPSE, as shown in  § 3; the last two are templates used in exploring the impact of removing hints about hidden space features, as shown in § 4.5.
@@ Appendix A Task Prompts Involved
Appendix B A Case Study of Numeric Suffixes
During our early experiments, we discover an intriguing phenomenon. Without any specific hints, the attacker LLM tend to actively generate some numeric suffixes. These numbers, when append directly to the given malicious query and input into the LLM, significantly increase the likelihood of the LLM continue writing extensive and diverse harmful content. Note that we explore the LLM itself here, i.e., the query is entered without using any other dialog template. However, when we replace these numbers with random numbers, the success rate of inducing harmful content drastically decrease. This observation suggests that LLMs may indeed possess and share certain special knowledge or memory, which can be exploited to induce harmful behaviors.
@@ Appendix B A Case Study of Numeric Suffixes
We make two further observations, with some empirical experiments on LLaMA2-7B-Chat. Firstly, we find that a numerical suffix may be able to activate certain specific memories in LLM. For example, as shown in  Table 6, adding a suffix “7.3” to Case 1 induces LLM to generate a sentence similar to a log mark and adding a suffix “5544.2” to Case 2 induces LLM to generate the address of a complementary pdf. Although we are currently unable to verify the authenticity of these logs or addresses, this responds to the possibility that LLM has seen similar content in the training data. Secondly, we try to analyze why adding a number causes LLM to behave differently. More specifically, we randomly select some malicious queries, enumerated tens of thousands of numeric suffixes for each query, and verify whether they induced harmful content. We visualize the embeddings of the prompts that successfully induced harmful content along with those that failed, wich PCA. Figure 4 shows a clear difference in the distribution of embeddings for failed prompts (red dots) and successful prompts (blue dots), which enlightens us that some numerical suffixes greatly perturb the hidden space features of the original malicious queries, thus destroying the robustness of the safety alignment.
@@
Appendix C A Study on Longitudinal Versions of GPT-3.5-Turbo
@@
enhancing linguistic capabilities, may have compromised some aspects of robustness and security Liu et al. (2023b). To explore this, we conduct a study on three versions of the GPT-3.5-Turbo model, specifically the 2023-0613, 2023-1106, and 2024-0125 releases, with ECLIPSE. As illustrated in  Figure 5, our results also corroborate this viewpoint. With updates across three versions of GPT-3.5-Turbo, the ASR increases progressively from 0.37 to 0.81, and then to 0.97, indicating a growing likelihood of producing harmful content.